A solenoid, 1.2meters long is uniformly wound with 800 turns.A search coil of mean diameter 30millimeter wound with 50 turns, is place at the center of the solenoid and connected to a ballastic galvanometer.The total resistance of the search coil circuit is 2000ohms. When a current of 5ampere through the solenoid is reversed, the deflection on the galvanometer scale is 85 division. Taking free space permeability 4π×10^-7Hm. Determine the ballastic constant of the galvanometer in coulombs per unit deflection?

1
Expert's answer
2021-07-19T03:15:14-0400

"B_o=\frac{4\pi *10^{-7}*800*5}{1.2}=4.18*10^{-3}T\\
r= \frac{D}{2}=\frac{3}{2}=1.5 mm= 1.5*10^{-3}m\\
q=\frac{2(4.18*10^{-3}(50)*(\pi)(1.5*10^{-3})^2 cos 0}{2000}=1.47**10^{-9}C\\
k= \frac{1.47*10^{-9}}{85}=1.73*10^{-11} C/cm"